Gunnar Corp. uses no debt. The weighted average cost of capital is 8.1 percent. The current market value of the equity is $40Kolby Corp. is comparing two different capital structures. Plan I would result in 9,000 shares of stock and $80,000 in debt.
Answer #1

1) Here, Gunnar Corp uses no Debt.

Given data WACC = 8.1% , Market value of Equity = $ 40 Million, Tax rate = 22%.

Return on Equity equals WACC as there is no Debt. Hence Re = 8.1%

Net Income = Return on Equity * Market value of Equity

= 8.1% * 40,000,000

= $ 3,240,000

=> Net Income = $ 3,240,000

EBIT = Net Income /(1-Tax rate)

= 3,240,000/(1-0.22)

= 3,240,000/0.78

= $ 4,153,846.15

=> EBIT of  Gunnar Corp is $ 4,153,846.15

2) Calculate under M & M proposition without taxes

Plan 1 - 9000 shares of stock and Debt = $ 80,000

Interest rate on debt = 8%, EBIT = $ 50,000, All Equity plan of 12000 shares.

Price per share of equity = value of shares repurchased / Number of shares repurchased

= 80000 / (12000-9000)

= $ 26.67

Hence under plan 1, Price per share of equity = $ 26.67

Plan 2- 7500 shares of stock and Debt = $ 120,000

Interest rate on debt = 8%, EBIT = $ 50,000, All Equity plan of 12000 shares.

Price per share of equity = value of shares repurchased / Number of shares repurchased

= 120000 / (12000-7500)

= $ 26.67

Hence under plan 2, Price per share of equity = $ 26.67